Breakdown of a cartel agreement 3. Breakdown of a cartel agreement Aa Aa E Imagine that there is a town in which only two of the residents, Stelios and Sondr After Stelios implements his new plan, the price of water and Stelioss production levels, Stelioss profit becomes _ per lit
Answer #1

Zitres demanded Tқ ( proE) 20 Чор Oo8 (обо 1200 І тро 2006 2700 З 200 35 3600 - топ 3500 32 00 2106 3 1406 1 боб 16об 2офо 20untel Acort Profit max puce is $3 . Total output is 1200 litres sebess fofit = 3600/2 = $1800 v sondras pussit = 3600/2 = $1Duuruus (313-152) S FALSE. Both the producers that the output effect was puie effect assumed greater than because autput effe